# Evaluating Sound Similarity Metrics for Differentiable, Iterative Sound-Matching

### v1

@@ II-E Iterative Sound Matching
Horner et al. [8] Non-diff FM STFT (McAulay-Quatieri) Direct Replicate In 1991 Iter
Mitchel et al. [12] Non-diff FM Spectral Relative Evolutionary Replicate In (Contrived) 2007 Iter
Yee-King et al. [17] Non-diff VST P-Loss (for Supervised), Spectral (MFCC) Supervised & Evolutionary Replicate In 2018 Both
Esling et al. [10] Non-diff VST Spec MSS/SC P-loss Supervised & Modeling Replicate In* & Out 2019 1-shot
@@
Salimi et al. [9] Non-diff STFT and Envelope Supervised Imitate Out 2020 1-shot
@@
Vahidi et al. [15] Differentiable FM JTFS/MSS Direct (grad desc.) Replicate In 2023 Iter
@@
Uzrad et al. [27] DDSP chains Synth. Chain and P-loss Supervised Replicate In & Out 2024 1-shot
III Past Works
Table I contains an overview of relevant literature, covering important works in sound-matching. Earlier works in this domain focused on FM synthesis and genetic algorithms (or direct optimization methods)[16, 8, 12]. Along with the exponential increase in computation power, later works utilized a wider range of approaches to parameter search, estimation, and audio synthesis; particularly with supervised learning and neural networks (NN) [17, 11, 10].
III-A Earlier Works
FM synthesizers are simple to implement, yet very expressive [46]. Perhaps the earliest foundational work in sound-matching is the analytical approach by Justice [16] toward the decomposition and recreation of sounds with FM synthesis using one carrier, one modulator, and the corresponding envelopes:
$f(t)=I_{c}(t)\cos(\omega_{c}t+I_{m}(t)\cos(\omega_{m}t))$
Where $f(t)$ is the output signal, $I_{c}(t)$ is the carrier amplitude, $\omega_{c}$ is the carrier frequency, and $I_{m}(t)$ and $\omega_{m}$ are the envelope and frequency for the modulator [16].
The sound-matching works that followed used other heuristics for finding the synthesizer parameters, yet the simple FM synthesizer structure remained unchanged. Horner et al. [8] used GAs for re-synthesis of sounds with an FM synthesizer using one modulator and one carrier oscillator, and the McAulay-Quatieri method [47] for measuring loss.
Introduction of more complex models of synthesis such as wavetables [48] and physical modeling [32] rendered interesting results, but led to further questions about the nature and goals of sound-matching. Mitchell and Creasy noted the difficulty in distinguishing between “inefficiency of the optimization engine” and “synthesizer limitations” as the cause of failure or success in sound-matching [12]. They offered a contrived methodology for finding the best evolutionary method for sound-matching using their FM synthesizer. The methodology postulates that the best search heuristic for in-domain search (where an exact target exists, and a wide range of targets can be produced by sampling from different points in the parameter space) would also be the best-performing heuristic for out-of-domain search on a dataset of muted trumpet tones [49]; however, testing was not extensive and yielded some contradictory results. As they noted, modifications to the synthesizer, loss function, and sound domain would lead to a problem with an entirely new search space [12].
III-B Introduction of Deep-Learning
@@
Recent years have seen more works in sound-matching using machine learning techniques. In 2018, Yee-King et al. rendered 60,000 audio-parameter pairs from the Dexed VST synthesizer  [^1] https://asb2m10.github.io/dexed/, and showed that NNs trained on this dataset can outperform GA and hill-climber (HC) [50] methods in rendering speed, with slight improvements in MFCC error (used as an objective performance test). The speed improvement appears trivial, considering the iterative nature of GAs when compared to offline training of supervised models. For GAs and HC optimizer, MFCCs were used as a measure of performance as well as a loss function. For training the networks, P-loss was used, since differentiable MFCCs were not possible in their pipeline. Importantly, informal hearing tests revealed that the performance of even the best NN model was unsatisfactory, possibly due to the complex nature of the synthesizer, which features 155 parameters.
Esling et al. noted the issue of using P-Loss in sound-matching [10], considering the non-linear mapping between parameters of a synth and the audio output. In their approach, dubbed as “flowsynth”, they used a large parameter-to-audio dataset for a commercial VST synthesizer. In flowsynth, audio is encoded to an auditory space, using a NN autoencoding approach [51]. The encodings in the auditory space can then be converted to the parameter space, but only for that particular synthesizer. Using P-Loss and Spectral Convergence (Frobenius norm of the spectrograms) as metrics, flowsynth outperformed other NN approaches that did not use an auditory to parameter space conversion, but formal hearing tests were not conducted.
@@
Masuda et al. also applied supervised learning to synthesizer parameter estimation [13]. Their work highlights the issue of non-linearity in parameter-to-synthesizer outputs and out-of-domain search. This work uses a differentiable subtractive synthesizer (two oscillators and an LP filter). A NN model was pre-trained using an in-domain dataset of randomly selected parameters and P-Loss. After training, the model was fine-tuned using 20,000 out-of-domain sounds from the NSynth dataset [52] and multi-scale spectrogram loss [11]. This approach proved more effective—that is, lower multi-level spectral difference in out-of-domain tests—than baseline models, which were either not exposed to out-of-domain sounds or trained exclusively with P-Loss. Subjective hearing tests were conducted, showing a preference for the fine-tuned model [13]. Masuda et al. later extended this work with semi-supervised learning, highlighting significant gaps in in-domain and out-of-domain performance [14].
@@
Rather than focusing on a particular implementation, Shier et al. presented Spieglib, a library for implementation of sound-matching pipelines [5]. This library provides different choices for DNNs, GAs, synthesizers, and feature extractors. Shier et al. presented an experiment with a similar setup to Yee-King et al. [17], however, they found a genetic algorithm to be the best performing.
@@
Differentiable loss functions that use spectrogram differences can be computationally expensive. To mitigate this, Han et al. [18] introduced “perceptual-neural-physical loss” (PNP). PNP is an approximation of loss functions; specifically, a loss function that uses the L2 norm of the difference between the features of two sounds, or $||\phi(t)-\phi(x)||^{2}_{2}$, where $\phi$ could be a spectrogram or JTFS function. PNP loss functions are fast and differentiable, but require training and parameter estimation. A Riemannian metric M needs to be calculated for the minimization of locally linear approximation of the “true” spectral loss function [18].
$\|\phi(t)-\phi(x)\|_{2}^{2}=\langle\tilde{\theta}-\theta|\mathbf{M}(\theta)|\tilde{\theta}-\theta\rangle+O(\|\tilde{\theta}-\theta\|_{2}^{3}).$
@@
During the training phase, this metric is calculated alongside the neural parameter estimator. After training, PNP can be used as a fast approximation of computationally demanding loss functions such as JTFS [18, 19], paired with FM synthesizers and a differentiable physical model of a drum head [53].
@@
III-C What is Lacking In the Field
Having looked at the current available literature, we note four major areas of weakness in sound-matching. We target the first two issues in this work, while the latter two issues are left for future work.
@@ III-C What is Lacking In the Field
    Loss Selection: Is there a best performing loss function, or is the selection of the loss function dependent on factors such as the sound domain, synthesizer capability and the desired characteristics of the output?
@@
    Synthesis Selection: There is a lack of diversity in the synthesis methods used in sound-matching. The use of different DSP functions for iterative sound-matching and their effects on the outcome or the interactions with the loss function have not been tested.
@@
    Periodic Loss: The sinusoidal nature of loss function landscapes is a problem that appears frequently. In differentiable settings, this problem causes gradient descent updates to local minima.
@@
    Out-Domain Generation: Sound-matching with out-of-domain sounds is an unexplored area, yet a necessary one for practical applications for sound designers.
IV Programs, Methodology, and Results
Towards addressing the problems of Loss Selection and Synthesis Selection, we compare the performance of iterative sound-matching experiments, where four differentiable loss functions are applied to four differentiable synthesizers. What we mean by “performance” is the similarity of the final output to the target, as measured by P-Loss, MSS, and—most importantly—manual listening tests.
We hypothesize that the relative performance of loss functions for sound-matching is influenced by the method of synthesis. If true, this would strengthen the idea that the selection of the loss function and synthesis method is not an optimization problem, but a subjective choice depending on the needs of the sound designer.
IV-A Experiment Methodology
We test four different loss functions and compare their success in sound-matching using four different synthesizers. We verify the success of the sound-matching process using two automatic measures of sound similarity (MSS and P-Loss) and a manual blinded hearing test conducted by two of the authors where Likert scores [54] of 1-5 are assigned to a subset of final outputs and the corresponding targets.
An experiment is one iteration of sound-matching given a loss function, target sound, and synthesizer program. When the final iteration is complete, a score is given (automatically with an algorithm or with manual listening) to an experiment that measures its success in replicating the target sound. Consider the following simplified scenario: we have two loss functions, $L_{1}$ and $L_{2}$, and a single synthesizer program $P$ with a target parameter set $\theta^{*}$ where the target sound is $t$, or $P(\theta^{*})=t$. We run a large number (in this case, 300 for each loss function) of experiments where we randomly set the initial ($\theta_{0}$) and target parameters ($\theta^{*}$) of $P$ and run an iterative sound-matching loop for each loss function, with maximum number of iterations set to 200. When the loop is terminated, we measure the similarity of the final output sound $x$ (where $P(\theta_{199})=x$) to $t$, or if using P-Loss we compare $\theta_{199}$ to $\theta^{*}$. Having 300 experiments gives us a distribution of losses to determine whether $L_{1}$ or $L_{2}$ led to better performance. This simplified example leads us to our methodology which has four loss functions, four synthesizer programs, and three different evaluation methods of scoring sound similarity.
IV-A1 Automatic and Manual Evaluation
The two automatic evaluation methods used here are P-Loss and MSS, as described in Section II-B. For P-Loss, the parameters are normalized between 0-1 based on the valid ranges defined in the Faust program, and the L1 distance between the normalized parameters is calculated. MSS is computed using a hop size of $100$ samples, and FFT window lengths of $(512,1024,2048,4096)$.
Automatic sound evaluation methods are often unreliable, as their alignment with the subjective judgments of human listeners is uncertain. Therefore, listening tests are conducted by randomly sampling 40 experiment results for each program and loss function pair. Two of the authors then assign blinded similarity scores to the outputs and targets using a 5-point Likert scale (1 = no similarity, 5 = near identical) [54]. Using Spearman’s rank correlation [55, 56], we found a very strong rate of agreement between the human listeners (see Section IV-E), a good indicator that the scores assigned by manual listeners can be treated as the ground truth. We then combine the ranks assigned by both authors, giving us 80 ranks per program and loss function pair.
IV-A2 Determining Best Performers
Figure 2 is a visualized summary of how the loss functions are ranked. We have four loss functions and four synthesizer programs. The maximum number of iterations set to 200, and 300 experiments for each loss/synthesizer combination. Each synthesizer program is paired with four loss functions and 300 experiments are conducted and evaluated automatically with P-Loss and MSS. Two of the authors assigned blinded Likert scores to 40 randomly selected sounds, where integers from 1 to 5 are assigned, with 1 meaning no similarity and 5 for near identical.
For each program and loss function pair, we have three distributions that can be used to rank the loss functions. Two with 300 automatically assigned similarities (P-Loss and MSS) and one with 80 Likert scores (combining the 40 ranks assigned by each author). For consistency and statistical robustness, the distributions are upsampled to 1000 values using bootstrapping [57]. Bootstrapping gives an estimation of the distribution for the mean performance of each experiment. $k$ (set to 1000) samples of $n$ values (set to 100% of the values) are taken with replacement from the empirical distribution (list of 300 performance values for MSS and P-Loss or 80 for manual rankings). The mean performance is calculated for each of the $k$ samples. The $k$ estimates of the mean performance give us a bootstrapped distribution for each loss function.
For each program, the loss functions are ranked using the non-parametric Scott-Knott (NPSK) test [58, 59], which ranks the loss functions from rank 1 (best, or highest distribution values) to maximum rank of 4 (worst, with lowest distribution values). NPSK may determine that multiple functions belong to the same rank, in which case the maximum rank would be less than 4. As mentioned before, lower P-Loss and MSS values indicate higher similarity, therefore (for consistency with manually assigned ranks) the reciprocal of their output is used ($x^{-1}$, if $x$ is the P-Loss or MSS value), where higher values indicate greater similarity.
[…]
IV-A3 Training Parameters and Gradient Calculations
IV-B Loss Function Implementation Details
IV-B1 STFT Losses
Due to their ubiquity and lower cost of gradient calculation, we use STFT as the basis of two of the loss functions. As discussed in Section II-D, the L1 or L2 distances are the most common methods of comparing different spectrograms, however, there are situations where this might lead to shortcomings. An example given by Vahidi et al. [15] is two chirplets (tones that increase in pitch) starting and ending at the same frequencies. If one of these tones is slightly shifted in time, then the spectrograms will no longer overlap at any point, despite their sonic similarity. In this example, we do not know if this is an issue of the spectrograms, or the L1 measure used for their comparison. In the computer vision field, Structural Similarity Index (SSIM) [61, 62] and Scale-Invariant Mean Squared Error (SIMSE) [63] have been used as alternatives that improved performance in various tasks relative to L1.
@@ IV-B2 JTFS Loss
IV-B3 Soft-DTW Loss
We use the soft-DTW function, which is differentiable and — depending on its parameters — not shift-invariant [65, 66, 67]. The loss function DTW_Envelope is the application of the soft-DTW function to the envelope of the two sounds being compared [3]. The envelope is calculated by creating the STFT spectrogram of a sound (the same process used in Section II-B2) and summing the values at each timestep.
IV-C The Synthesizers

### v2

@@ II-E Supervised Versus Direct Optimization
Horner et al. [6] Non-differentiable FM STFT (McAulay-Quatieri) GA Replicate In 1991 Iter
Mitchel et al. [10] Non-differentiable FM Spectral Relative Evolutionary Replicate In (Contrived) 2007 Iter
Yee-King et al. [30] Non-differentiable VST P-Loss (for Supervised), Spectral (MFCC) Supervised & Evolutionary Replicate In 2018 Both
Esling et al. [8] Non-differentiable VST Spec MSS/SC P-Loss Supervised & Modeling Replicate In* & Out 2019 1-shot
@@
Salimi et al. [7] DSP STFT and Envelope Supervised Imitate Out 2020 1-shot
@@
Vahidi et al. [14] Differentiable FM JTFS/MSS Gradient desc. Replicate In 2023 Iter
Barkan et al. [40] Diff. Synthesizer Proxy P-Loss + L1 Spec (weighted) Supervised (init.), Semi-Supervised Replicate In 2023 Both
@@
Uzrad et al. [15] DDSP Synth. Chain and P-Loss Supervised Replicate In & Out 2024 1-shot
Cherep et al. [41] DDSP CLAP Evolutionary Imitate Out 2024 Iter
II-F Historical Framing of Sound-Matching
Given the attributes discussed above, Table I summarizes relevant literature. Here we expand with a historical analysis of past sound-matching work.
Perhaps the earliest foundational study is Justice [42], who analytically decomposed and recreated sounds using a simple FM synthesizer (see Section II-B). Subsequent work retained this FM structure but explored new heuristics for parameter search. For example, Horner et al. [6] applied genetic algorithms (GAs) to resynthesize sounds with one modulator and one carrier, measuring similarity via the McAulay–Quatieri method [43].
Later studies introduced more complex synthesis models, such as wavetables [44] and physical modeling [21]. Mitchell and Creasy [10] highlighted the difficulty of disentangling synthesizer limitations from optimization inefficiency. They proposed a contrived methodology in which the best heuristic for in-domain FM resynthesis should also generalize to out-of-domain targets (e.g., muted trumpet tones [45]). However, limited testing produced contradictory results, and they concluded that changes to the synthesizer, loss, or sound domain effectively redefine the search space [10].
@@ II-F Historical Framing of Sound-Matching
Recent years have seen more works in sound-matching using supervised machine learning techniques. In 2018, Yee-King et al. rendered 60,000 audio-parameter pairs from the Dexed VST synthesizer  [^1] https://asb2m10.github.io/dexed/, and showed that NNs trained on this dataset can outperform GA and hill-climber (HC) [46] methods in rendering speed, with slight improvements in MFCC error (used as an objective performance test). The speed improvement appears trivial, considering the iterative nature of GAs when compared to offline training of supervised models. For GAs and HC optimizer, MFCCs were used as a measure of performance as well as a loss function. For training the networks, P-Loss was used, since differentiable MFCCs were not possible in their pipeline. Importantly, informal listening tests found the results unsatisfactory, likely due to the synthesizer’s 155-parameter complexity.
@@
Masuda et al. also applied supervised learning [12]. Their work highlights the issue of non-linearity in parameter-to-synthesizer outputs and out-of-domain search. This work uses a differentiable subtractive synthesizer (two oscillators and an LP filter) alongside a NN model pre-trained with P-Loss on an in-domain dataset of randomly selected parameters. After training, the model was fine-tuned using 20,000 out-of-domain sounds from the NSynth dataset [47] and multi-scale spectrogram loss [9]. This approach proved more effective—i.e., achieved lower multi-level spectral difference in out-of-domain tests—than baseline models, which were either not exposed to out-of-domain sounds or trained exclusively with P-Loss. Subjective hearing tests were conducted, showing a preference for the fine-tuned model [12]. Masuda et al. later extended this work with semi-supervised learning, highlighting significant gaps in in-domain and out-of-domain performance [13].
@@
Rather than focusing on a particular implementation, Shier et al. presented Spieglib, a library for implementation of sound-matching pipelines [11]. This library provides different choices for DNNs, GAs, synthesizers, and feature extractors. Shier et al. presented an experiment with a similar setup to Yee-King et al. [30], however, they found a genetic algorithm as the best performing.
@@
Differentiable loss functions that use spectrogram differences can be computationally expensive. To mitigate this, Han et al. [16] introduced “perceptual-neural-physical loss” (PNP). PNP is an approximation of loss functions; specifically, a loss function that uses the L2 norm of the difference between the features of two sounds, or $\phi(t)-\phi(x)||^{2}_{2}$, where $\phi$ could be a spectrogram or JTFS function. PNP loss functions are fast and differentiable, but require training and parameter estimation. A Riemannian metric M needs to be calculated for the minimization of locally linear approximation of the “true” spectral loss function [16].
$\|\varphi(x_{0})-\varphi(x_{\theta})\|_{2}^{2}=\langle\tilde{\theta}-\theta\,|\,M(\theta)\,|\,\tilde{\theta}-\theta\rangle+O(\|\tilde{\theta}-\theta\|_{3}^{2}).$
@@
This metric is calculated alongside the neural parameter estimator. Once trained, PNP enabled the fast differentiable optimization of computationally expensive loss functions such as JTFS with FM and physical models [16, 22].
A neural approximation for another part of the supervised sound-matching chain—this time the synthesizer—was proposed by Barkan et al. [40]. As they noted, without a differentiable synthesizer, “model-based” (or what we call supervised) approaches cannot directly compare $x_{\theta}$ to $x_{0}$, often opting for P-Loss, which may not correctly map the parameters of sound to the output audio [8, 16, 13]. This approach used a model for mapping sounds to parameters, another for approximating the synthesizer, and a loss function which combines P-Loss with STFT differences. This “Inversynth II” (IS2) approach yielded significant improvements to previous works which did not use the STFT approximations for loss [8, 48]. Barkan et al. then attempted to improve the IS2 model with Inference-Time Fine-tuning (ITF). For ITF, the synthesizer approximation is frozen and the encoder is iteratively fine-tuned for a particular sample. However, ITF often degraded performance, possibly due to proxy–synthesizer mismatch or overfitting.
@@
Audio embeddings can also be used as a similarity metric. In a recent work, Cherep et al. used latent representations from the CLAP model [49] along with a differentiable synthesizer [50] to create creative interpretations of sound effects. In their approach, a desired sound-effect is described in text and embedded using CLAP, followed by iterative updates from a gradient-free optimizer [51] to the synthesizer’s parameters to minimize embedding differences. Based on manual hearing tests, this approach did not produce the “correct” sounds more frequently than previous works [52]. However, it did yield better scores for “artistic-interpretation” (what we referred to as “imitation”). It should be noted that audio embedding models are often trained using simpler loss functions—particularly those which utilize Fourier transforms—therefore improving such simpler loss functions is likely to improve the embedding models.
II-G What is Lacking In the Field
Having looked at the current literature, we identify four major areas of weakness in sound-matching. We target the first two issues in this work, while the latter two are left for future work.
@@ II-G What is Lacking In the Field
    Loss Selection: It remains unclear whether there exists a universally best-performing loss function, or whether performance depends on factors such as the sound domain, synthesizer architecture, and desired output characteristics. Existing studies typically evaluate losses only with simple synthesis setups [14], leaving open questions about their generality.
@@
    Synthesis Selection: There is a lack of diversity in the synthesis methods used in sound-matching. The effects of using different isolated DSP functions in iterative sound-matching and how these interact with loss functions remains largely untested.
@@
    Loss Landscape Navigation: Loss function landscapes are not easy to navigate, as there can be many local minima, or large flat areas [5, 14]. In differentiable settings, this can cause gradient descent updates to not reach the global minima.
@@
    Out-of-Domain Generation: Sound-matching with out-of-domain sounds is an under-explored area, yet a necessary one for practical applications for sound designers.
II-H What Approach Should We Take?
We argue that isolated benchmarking is a necessary precursor to building robust, generalizable sound-matching systems. To address the problems of Loss Selection and Synthesis Selection, we adopt a methodology with three key characteristics that have not previously appeared together.
  1. 1.
    Use simple differentiable synthesizers built from common DSP functions (oscillators, filters, envelopes) to address the Synthesis Selection problem. Restriction to small-scale, low-parameter models avoids the confounding effects of neural proxies or embedding models, enabling direct analysis of how synthesis method interacts with loss functions.
  2. 2.
    Evaluate experiments under multiple loss functions to address the Loss Selection problem. Prior work has typically tested losses in isolation, making direct comparisons of loss functions difficult [14, 16, 15]. Particularly rare are comparisons which are done in the context of different methods of synthesis.
  3. 3.
    Optimize synthesizer parameters directly with gradient descent, without neural approximations. This complements prior proxy-based methods by providing the missing “low-level experiments” that clarify fundamental interactions between DSP functions and similarity measures.
Together, these design choices supply the controlled experimental evidence missing from the literature, clarifying how losses and synthesis methods interact in iterative sound-matching, and offering insights likely to generalize to more complex domains.
III Experimental Setup and Results
At a glance, our methodology is to conduct controlled, low-level differentiable experiments without the use of proxy networks. We pair four differentiable losses with four differentiable DSP-based synthesizers: BP-Noise (band-pass noise with LP/HP filters), Add-SineSaw (additive sine and saw oscillators), Noise-AM (noise modulated by an LFO), and SineSaw-AM (sine-modulated saw oscillator). We optimize parameters directly via gradients in order to isolate loss–synthesizer interactions. This controlled setup systematically evaluates the performance of iterative sound-matching pipelines, where “performance” is defined as the similarity of the synthesized output to a target. Similarity is assessed with P-Loss, MSS, and—most importantly—manual listening tests. This design allows us to directly test our central hypothesis: that the effectiveness of a loss function depends on the synthesis method, and that no universal “best” similarity measure exists.
Here we discuss the implementation of the four loss functions (Section III-A) across four differentiable synthesizers (Section III-B). For each loss–synthesizer pair, we run a large number of trials and evaluate their outcomes automatically and manually (Section III-C). “Experiment” in this context refers to one complete iterative sound-matching run, from random parameter initialization through iterative gradient-based optimization until termination. Scores from these experiments are then aggregated and compared to determine best-performing losses and synth–loss pairings.
III-A Loss Function Implementation Details
III-A1 STFT Losses
Due to their ubiquity and lower cost of gradient calculation, we use STFT as the basis of two loss functions. L1 or L2 distances are the most common methods of comparing spectrograms [5, 28]. However, these measures can be overly sensitive to global gain or minor misalignments, potentially overstating perceptual differences. To address this, we also test Scale-Invariant Mean Squared Error (SIMSE) as an alternative to L1 for comparing STFT spectrograms.
@@ III-A2 JTFS Loss
III-A3 Soft-DTW Loss
We use the soft-DTW function, which is differentiable and—depending on its parameters—not shift-invariant [54, 55, 56]. The loss function DTW_Envelope is the application of the soft-DTW function to the envelope of the two sounds being compared [1]. This loss uses the similarity of amplitude modulation patterns, which are perceptually important in many sound-design scenarios. The envelope is calculated by creating the STFT spectrogram of a sound (the same process used in Section II-C2) and summing the values at each timestep.
III-B The Synthesizers